OpenStudy (anonymous):

State the slope and y-intercept for the graph of 3x + y = -2

OpenStudy (lgbasallote):

rewrite as y = -3x - 2 can you find slope and y-intercept now?

OpenStudy (anonymous):

y = -3x -2 m = -3 and c (y intercept) = -2.

Parth (parthkohli):

y = mx + b m is the slope b is the y-intercept.

OpenStudy (kropot72):

The standard form of the equation of a straight line is y = mx +b where m is the slope and b is the y-intercept. We need to rearrange 3x + y = -2 to the standard form. First subtract 3x from both sides to get y = -3x - 2 Now can you find the values for m and b?

OpenStudy (anonymous):